OpenStudy (anonymous):

What is the value of r in the equation 3r – 6 = 18?

OpenStudy (anonymous):

add 6 to both sides to get 3r = 24 then divide by 3 to get r=24/6 which is r = 4

OpenStudy (anonymous):

i messed up it sould be r=24/3 which is r=8

OpenStudy (anonymous):

So what's the answer?

OpenStudy (anonymous):

r=8

OpenStudy (anonymous):

8

OpenStudy (anonymous):

yes 8